Scenario 1 You are the operations manager of a firm that uses the continuous inventory control system. Suppose the firm operates 50 weeks a year, 350 days, and has the following characteristics for its primary item: Demand = 500 units/week Price paid to supplier = $160/unit Ordering cost = $80/order Holding cost = $16/unit/year Lead time = 1.5 weeks Standard deviation in weekly demand = 36 units Use the information in Scenario 1. Which of the following changes would result in the biggest increase in economic order quantity?

A- The ordering cost increases by 20% B- The demand increases by 25% C- The holding cost increases by 30% D- None of the above results in an increase in EOQ
Answer #1

Answer: - Option"B" --- The Demand increases by 25%

Explanation: - According to given data

Formula to calculate economic order quantity = √2*D*S/H

Annual demand = 500*50 = 25000

For given scenario

EOQ = √2*25000*$80/$16 = 500 units

If Ordering cost incresases by 20 percent, Order cost = $96

So EOQ = √2*25000*$96/$16 = 548 units

If demand increases by 25% then, Demand = 31,250

EOQ = √2*31250*$80/$16 = 559 units

If holding cost increases 30% holding cost = $20.8

EOQ = √2*25000*$80/$20.8 = 439 units

From the analysis we can conclude that Option B that is increasing demand by 25% increases EOQ.